OpenStudy (anonymous):

What is the equation of the line passing through (-3, 6) and parallel to the line y = 4x - 2 in standard form?

OpenStudy (anonymous):

\[y - y1 = m (x - x1)\]

OpenStudy (snuggielad):

^^^^is correct \(\Huge \color {red}{ \mathbb {Welcome\:\:to\:\:OpenStudy!!}} \)